A 4.00 kg model rocket is launched, shooting 50.0 g of burned fuel from its exhaust at an
average velocity of 625 m/s. What is the velocity of the rocket after the fuel has burned?
(Ignore effects of gravity and air resistance.)
1
Expert's answer
2011-08-25T13:21:25-0400
Let m1 be the mass of the rocket, m2 - the mass of burned fuel.
V1 - velocity of the rocket after the fuel has burned,
V2 - average velocity of burned fuel.
According to the law of conservation of linear momentum:


0 = m1V1 - m2V2


V1 = V2*m2 / m1
V1 = 625*0.05/4 = 7.8125 m/s


So, the velocity of the rocket is 7.8 m/s.